OpenStudy (anonymous):

Which value of x makes the following equation true? 2(x - 1) = -3x + 38 A. 8 B. 9 C. 12 D. 11

OpenStudy (anonymous):

@Write0The0Right0For0Wrong0Doing @willsgirl16

OpenStudy (texaschic101):

2(x - 1) = -3x + 38 --- distribute through the parenthesis 2x - 2 = -3x + 38 --- add 3x to both sides 2x + 3x - 2 = 38 -- add 2 to both sides 2x + 3x = 38 + 2 -- combine like terms 5x = 40 -- divide both sides by 5 x = 8 any questions ?

OpenStudy (anonymous):

Nope, thank you

OpenStudy (texaschic101):

no problem :)